a log cabin is 24 years old and a friendship quilt is 6 years old. In how many years will the log cabin be three times as old as the friendship quilt?

Answer by fractalier(6550)

Call the number of years in the future, x.
Thus we can write
24 + x = 3(6 + x)
24 + x = 18 + 3x
6 = 2x
x = 3 years from now